# InterAnimate: Taming Region-aware Diffusion Model for Realistic Human Interaction Animation

### v1

@@ 4.2 ID Preserver
$\mathbf{V}^{\mathrm{o}}=\mathbf{V}^{\mathrm{o}}_{att}+\mathbf{V}^{\mathrm{o}}_{\mathrm{face}}.$ (9)
@@
Figure 4: Qualitative comparison vs SOTA methods.
4.3 Training Objective
Following the previous methods [3, 5], InterAnimate is constructed as a latent video diffusion model. The training objective is formulated by incorporating diffusion loss to guide the denoising process, along with orthogonality loss to maximize the diversity of the patterns learned by the interaction latents.
@@ 4.3 Training Objective
$\mathcal{L}_{\text{diff}}=\mathbb{E}_{z}\left[\|z-\hat{z}\|_{2}^{2}\cdot\mathrm{W}_{\mathrm{hand}}\cdot\mathrm{W}_{\mathrm{face}}\right],$ (10)
$\mathrm{W}_{\mathrm{r}}=\begin{cases}\lambda_{\mathrm{r}}&\text{if}\ \mathrm{M}_{\mathrm{r}}>0,\\ 1.0&\text{ortherwise},\end{cases}$ (11)
@@
Table 1: Quantitative comparisons of InterAnimate with the recent SOTA methods.
@@
FID $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ LISPIS $\downarrow$ L1 $\downarrow$ FID-VID $\downarrow$ FVD $\downarrow$
Disco [45] CVPR 2024 116.04 0.482 18.00 0.274 3.65E-04 126.66 1292.37
AnimateAnyone [14] CVPR 2024 112.67 0.392 13.08 0.512 2.32E-04 100.71 964.11
MagicAnimate [52] CVPR 2024 45.93 0.509 15.14 0.305 1.37E-04 33.15 446.19
MimicMotion [61] arXiv 2024 55.02 0.458 14.85 0.363 7.50E-05 36.87 498.77
DisPose [21] ICLR 2025 33.30 0.565 17.04 0.246 1.07E-04 21.22 338.40
StableAnimator [40] CVPR 2025 24.95 0.748 22.24 0.162 2.69E-05 11.80 176.20
@@
Orthogonality Loss. This loss is introduced to encourage the interaction latents to be orthogonal to each other. This property is particularly important in the context of interaction patterns learning, where the learnable latents should ideally capture diverse and non-redundant interaction information to facilitate high-quality human intertaction animation. The Algorithm 2 shows the computation of the orthogonality loss function $\mathcal{F}_{ortho}(.)$. The orthogonality loss is applied separately to two distinct types of latents: spatial latents $\mathbf{L}^{\mathrm{s~{}}}$ and temporal latents $\mathbf{L}^{\mathrm{t~{}}}$. For each type, the loss is computed to reduce the similarity between any pair of latent vectors within the set. The total orthogonality loss $\mathcal{L}_{\text{ortho}}$ is then the sum of the spatial and temporal components, as given by:
$\mathcal{L}_{\text{ortho}}=\mathcal{L}_{\mathrm{s}}+\mathcal{L}_{\mathrm{t}},$ (12)
@@
Algorithm 2 Orthogonality Loss Calculation
0:  interaction latents $\text{interaction latents }\mathbf{L}\in\mathbb{R}^{n\times d}$
1:  $S\leftarrow\mathbf{L}\cdot\mathbf{L}^{T}\in\mathbb{R}^{n\times n}$triangleright similarity matrix
2:  $I\leftarrow\text{eye}(n)\in\mathbb{R}^{n\times n}$triangleright identity matrix
3:  $mask\leftarrow(1-\text{eye}(n))\in\mathbb{R}^{n\times n}$triangleright mask for non-diagonal
4:  $S_{\text{mask}}\leftarrow S[mask]$triangleright non-diagonal elements of S
5:  $I_{\text{mask}}\leftarrow I[mask]$triangleright non-diagonal elements of I
6:  $\mathcal{L}\leftarrow\text{MSE}(S_{\text{mask}},I_{\text{mask}})$triangleright Mean Squared Error
7:  Return: $\mathcal{L}$
@@
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{diff}}+\beta\cdot\mathcal{L}_{\text{ortho}},$ (13)
@@
5 Experiments
5.1 Implementation Details
Baselines and test set. We compare with 6 SOTA human image animation models, including Disco [45], AnimateAnyone [14], MagicAnimate [52], MimicMotion [61], DisPose [21] and StableAnimator [40]. To ensure fairness, we use our InterHF dataset to train the baselines that provide training code. To evaluate the model performance in interaction animation, we construct a test set that includes 18 types of interactions, with 10 videos for each type, sourced from different individuals, for a total of 180 videos.
Evaluation metrics. To rigorously evaluate the quality of image generation, we employ frame-level assessment using a suite of standard metrics, including Fréchet Inception Distance (FID) [11], Structural Similarity Index Measure (SSIM) [46], Learned Perceptual Image Patch Similarity (LISPIS) [60], Peak Signal-to-Noise Ratio (PSNR) [13], and L1 distance. For video evaluation, we construct clips by concatenating every 16 consecutive frames, and compute Video Fréchet Inception Distance (FID-VID) [2] and Fréchet Video Distance (FVD) [41] to jointly assess visual fidelity and temporal coherence. We utilize the evaluation code from Disco  [45].
Training. We choose the Stable Video Diffusion [3] as our architecture, and adopt the pre-trained weights from the public stable video diffusion 1.1 image-to-video model. Besides, we leverage the sampling strategy of Mimicmotion  [61]. The training is an end-to-end process, and the trainable components include denoising U-Net, pose encoder, learnable interaction latents and ID preserver. We train our model on 8 NVIDIA A100 GPUs for 50K steps with image size 384 $\times$ 640, learning rate 1e-5, and a per-device batch size of 1. During training, a random frame from the video is selected as the reference, and a 16-frame clip is sampled as the target. In our experiments, the learnable interaction latents have a dimension of $d=512$, while both spatial and temporal latents consist of 512 vectors each. During the soft quantization process, we employ a temperature parameter $\tau=1.0$. We set $\alpha$ in $\mathrm{Mixer}$ to 0.5 in our experiments. The loss amplification factors are set to $\lambda_{\mathrm{hand}}=5.0$ and $\lambda_{\mathrm{face}}=2.0$, respectively. Additionally, the orthogonality loss factor is established at $\beta=0.0001$.
5.2 Qualitative Comparison
@@
5.3 Quantitative Comparison
The quantitative results are summarized in Tab. 1, where multiple metrics are reported for both image and video quality. In terms of image quality metrics, InterAnimate outperforms all other methods. InterAnimate achieves the lowest FID (23.22), indicating high visual fidelity and alignment with real data. InterAnimate also records the highest SSIM (0.785) and PSNR (23.36), reflecting strong preservation of fine details and structure, which is crucial for realistic hand-face interactions like finger articulation and facial muscle movement. Additionally, the lowest L1 error and LPIPS further highlight InterAnimate’s precision and minimal distortion in reconstructing complex interactions. For video quality assessment, InterAnimate achieves the lowest FID-VID (10.79) and FVD (122.90), demonstrating its strength in generating temporally coherent and realistic videos. These results highlight its effectiveness in capturing natural motion in hand-face interactions, such as the smooth transition of fingers touching or moving away from the face.
@@ 5.3 Quantitative Comparison
5.4 Ablation Study
The ablation study in Tab. 2 examines the impact of key components, including the Region-aware Interaction System (RIS), soft quantization (quantize), orthogonality loss (o-loss), and ID Preserver (ID). The results demonstrate that RIS is essential for interaction animation, as its removal leads to a substantial decline across all image and video metrics. This highlights its importance in capturing fine-grained interaction details and maintaining temporal coherence in hand-face interactions. Removing quantization from the region attention block also causes a sharp decline in performance, highlighting its indispensable role in the extraction of priors from interaction latents. After removing the orthogonality loss and ID Preserver, there is a slight decrease in some metrics, with improvements observed in L1 and FID-FVD. However, the FVD value increased significantly (from 122.90 to 174.02 and 165.03, respectively), indicating that without o-loss, the interaction latents fail to capture sufficient interactive information, resulting in degraded performance over time. Similarly, temporal ID inconsistency also leads to a degradation in the FVD metric.
Table 2: Ablation analysis of the key components.
@@ 5.4 Ablation Study
Figure 5: The visualization of the ablation.
@@ 6 Conclusion and Discussion
To improve interaction animation, we will make two attempts in the future. Firstly, we intend to expand our interaction video dataset with a wider variety of interaction types and more intricate interaction modes. Secondly, we will explore stronger perception ability of interactions.

### v2

@@ 4.2. ID Preserver
(9) $\mathbf{V}^{\mathrm{o}}=\mathbf{V}^{\mathrm{o}}_{att}+\mathbf{V}^{\mathrm{o}}_{\mathrm{face}}.$
@@
Figure 4. Qualitative comparison vs SOTA methods.
4.3. Training Objective
Following the previous methods (Blattmann et al., 2023a, c), InterAnimate is constructed as a latent video diffusion model. The training objective is formulated by incorporating diffusion loss to guide the denoising process, along with orthogonality loss to maximize the diversity of the patterns learned by the interaction latents.
@@ 4.3. Training Objective
(10) $\mathcal{L}_{\text{diff}}=\mathbb{E}_{z}\left[\|z-\hat{z}\|_{2}^{2}\cdot\mathrm{W}_{\mathrm{hand}}\cdot\mathrm{W}_{\mathrm{face}}\right],$
(11) $\mathrm{W}_{\mathrm{r}}=\begin{cases}\lambda_{\mathrm{r}}&\text{if}\ \mathrm{M}_{\mathrm{r}}>0,\\ 1.0&\text{ortherwise},\end{cases}$
@@
Table 1. Quantitative comparisons of InterAnimate with the recent SOTA methods.
@@
FID $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ L1 $\downarrow$ FID-VID $\downarrow$ FVD $\downarrow$
Disco (Wang et al., 2024a) CVPR 2024 116.04 0.482 18.00 0.274 3.65E-04 126.66 1292.37
AnimateAnyone (Hu, 2024a) CVPR 2024 112.67 0.392 13.08 0.512 2.32E-04 100.71 964.11
MagicAnimate (Xu et al., 2024c) CVPR 2024 45.93 0.509 15.14 0.305 1.37E-04 33.15 446.19
MimicMotion (Zhang et al., 2024) arXiv 2024 55.02 0.458 14.85 0.363 7.50E-05 36.87 498.77
DisPose (Li et al., 2025) ICLR 2025 33.30 0.565 17.04 0.246 1.07E-04 21.22 338.40
StableAnimator (Tu et al., 2025) CVPR 2025 24.95 0.748 22.24 0.162 2.69E-05 11.80 176.20
@@
Algorithm 2 Orthogonality Loss Calculation
[…]
4:SmaskS[$_{\text{mask}}\leftarrow S$[mask] {non-diagonal elements of S}
5:ImaskI[$_{\text{mask}}\leftarrow I$[mask] {non-diagonal elements of I}
[…]
7:Return: $\mathcal{L}$
Orthogonality Loss. This loss is introduced to encourage the interaction latents to be orthogonal to each other. This property is particularly important in the context of interaction patterns learning, where the learnable latents should ideally capture diverse and non-redundant interaction information to facilitate high-quality human intertaction animation. The Algorithm. 2 shows the computation of the orthogonality loss function $\mathcal{F}_{ortho}(.)$. The orthogonality loss is applied separately to two distinct types of latents: spatial latents $\mathbf{L}^{\mathrm{s~}}$ and temporal latents $\mathbf{L}^{\mathrm{t~}}$. For each type, the loss is computed to reduce the similarity between any pair of latent vectors within the set. The total orthogonality loss $\mathcal{L}_{\text{ortho}}$ is then the sum of the spatial and temporal components, as given by:
(12) $\mathcal{L}_{\text{ortho}}=\mathcal{L}_{\mathrm{s}}+\mathcal{L}_{\mathrm{t}},$
@@
(13) $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{diff}}+\beta\cdot\mathcal{L}_{\text{ortho}},$
@@
5. Experiments
5.1. Implementation Details
Baselines and test set. We compare with 6 SOTA human image animation models, including Disco (Wang et al., 2024a), AnimateAnyone (Hu, 2024a), MagicAnimate (Xu et al., 2024c), MimicMotion (Zhang et al., 2024), DisPose (Li et al., 2025) and StableAnimator (Tu et al., 2025). To ensure fairness, we use our InterHF dataset to train the baselines that provide training code. To evaluate the model performance in interaction animation, we construct a test set that includes 18 types of interactions, with 10 videos for each type, sourced from different individuals, for a total of 180 videos.
Evaluation metrics. To rigorously evaluate the quality of image generation, we employ frame-level assessment using a suite of standard metrics, including Fréchet Inception Distance (FID) (Heusel et al., 2017), Structural Similarity Index Measure (SSIM) (Wang et al., 2004), Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al., 2018), Peak Signal-to-Noise Ratio (PSNR) (Horé and Ziou, 2010), and L1 distance. For video evaluation, we construct clips by concatenating every 16 consecutive frames, and compute Video Fréchet Inception Distance (FID-VID) (Balaji et al., 2019) and Fréchet Video Distance (FVD) (Unterthiner et al., 2018) to jointly assess visual fidelity and temporal coherence. We utilize the evaluation code from Disco  (Wang et al., 2024a).
Training. We choose the Stable Video Diffusion (Blattmann et al., 2023a) as our architecture, and adopt the pre-trained weights from the public stable video diffusion 1.1 image-to-video model. Besides, we leverage the sampling strategy of Mimicmotion  (Zhang et al., 2024). The training is an end-to-end process, and the trainable components include denoising U-Net, pose encoder, learnable interaction latents and ID preserver. We train our model on 8 NVIDIA A100 GPUs for 50K steps with image size 384 $\times$ 640, learning rate 1e-5, and a per-device batch size of 1. During training, a random frame from the video is selected as the reference, and a 16-frame clip is sampled as the target. In our experiments, the learnable interaction latents have a dimension of $d=512$, while both spatial and temporal latents consist of 512 vectors each. During the soft quantization process, we employ a temperature parameter $\tau=1.0$. We set $\alpha$ in $\mathrm{Mixer}$ to 0.5 in our experiments. The loss amplification factors are set to $\lambda_{\mathrm{hand}}=5.0$ and $\lambda_{\mathrm{face}}=2.0$, respectively. Additionally, the orthogonality loss factor is established at $\beta=0.0001$.
5.2. Qualitative Comparison
@@
5.3. Quantitative Comparison
The quantitative results are summarized in Table 1, where multiple metrics are reported for both image and video quality. In terms of image quality metrics, InterAnimate outperforms all other methods. InterAnimate achieves the lowest FID (23.22), indicating high visual fidelity and alignment with real data. InterAnimate also records the highest SSIM (0.785) and PSNR (23.36), reflecting strong preservation of fine details and structure, which is crucial for realistic hand-face interactions like finger articulation and facial muscle movement. Additionally, the lowest L1 error and LPIPS further highlight InterAnimate’s precision and minimal distortion in reconstructing complex interactions. For video quality assessment, InterAnimate achieves the lowest FID-VID (10.79) and FVD (122.90), demonstrating its strength in generating temporally coherent and realistic videos. These results highlight its effectiveness in capturing natural motion in hand-face interactions, such as the smooth transition of fingers touching or moving away from the face.
@@ 5.3. Quantitative Comparison
5.4. Ablation Study
The ablation study in Table 2 examines the impact of key components, including the Region-aware Interaction System (RIS), soft quantization (quantize), orthogonality loss (o-loss), and ID Preserver (ID). The results demonstrate that RIS is essential for interaction animation, as its removal leads to a substantial decline across all image and video metrics. This highlights its importance in capturing fine-grained interaction details and maintaining temporal coherence in hand-face interactions. Removing quantization from the region attention block also causes a sharp decline in performance, highlighting its indispensable role in the extraction of priors from interaction latents. After removing the orthogonality loss and ID Preserver, there is a slight decrease in some metrics, with improvements observed in L1 and FID-FVD. However, the FVD value increased significantly (from 122.90 to 174.02 and 165.03, respectively), indicating that without o-loss, the interaction latents fail to capture sufficient interactive information, resulting in degraded performance over time. Similarly, temporal ID inconsistency also leads to a degradation in the FVD metric.
Table 2. Ablation analysis of the key components.
@@ 5.4. Ablation Study
Figure 5. The visualization of the ablation.
@@ 6. Conclusion and Discussion
To improve the interaction animation further, we will make two attempts in the future. Firstly, we intend to expand our interaction video dataset with a wider variety of interaction types and more intricate interaction modes. Secondly, we will develop models that are stronger in perception of interactions.
Acknowledgements.
This work was supported by Ant Group Research Intern Program and the Pioneer R&D Program of Zhejiang Province(No. 2024C01024). This work was also partially supported by Shenzhen Key Laboratory of next generation interactive media innovative technology(No: ZDSYS20210623092001004).